OpenStudy (anonymous):

f(t) = 1000e^0.2. Find f^-1(p)

OpenStudy (anonymous):

I know this is a log problem

OpenStudy (zzr0ck3r):

shuold there be a t in there somewhere?

OpenStudy (anonymous):

oh, yeah

OpenStudy (zzr0ck3r):

else it does not make sense

OpenStudy (anonymous):

1000e^0.2t

OpenStudy (zzr0ck3r):

so f(x)= 1000e^(0.2t) y = 1000e^(0.2t) we need to find the inverse y/1000 = e^(0.2t) ln(y/1000) = ln(e^(0.2t)) ln(y/1000)=0.2tln(e) = 0.2t so t = ln(y/1000)/0.2

OpenStudy (anonymous):

so when I solve the problem, do I substitute p in for y?

OpenStudy (anonymous):

since it is f^-1(p)

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

before just used x and y

OpenStudy (zzr0ck3r):

yes